OpenStudy (sbuck98):

Someone help me answer these 4 problems, I will award a medal, and coins.

OpenStudy (sbuck98):

1.) Slope (m)= change in y over change in x. M= y2-41 over x2-x1 show your work on how you found the slop 2.) Find the equation using the point slope formula. Point slope formula is y-y1=m(x-x1 3.) .Give a example of a line parallel to your answer from question #2. 4.) Give a example of a line perpendicular to your answer from #2

OpenStudy (anonymous):

To clarify, is M y^2 -41, or just y_2?

OpenStudy (sbuck98):

its just the variable Y2

OpenStudy (sbuck98):

@armyengineer225

OpenStudy (anonymous):

Well, the slope of an equation is just the derivative, or the change of y over x. However, it's slightly confusing because to equation variable is uppercase M.

OpenStudy (sbuck98):

its suppose to be lowercase. I have no idea how to do this, I just started this. @armyengineer225

OpenStudy (sbuck98):

@dan815

OpenStudy (sbuck98):

@FibonacciChick666 @haleyelizabeth2017

OpenStudy (fibonaccichick666):

so your questions make like zero sense to me at the moment

OpenStudy (fibonaccichick666):

can you type the exact wording and questions or screenshot or something?